OpenStudy (anonymous):

Find the quotient. (6x2 - x - 40) ÷ (5 + 2x)

OpenStudy (anonymous):

6x^2 - x - 40 = 6x^2 + 15x - 16x - 40 = 3x(2x + 5) - 8(2x + 5) = (3x - 8)(2x + 5). Thus, (6x^2 - x - 40)/(5 + 2x) = [(3x - 8)(2x + 5)]/(2x + 5) = 3x - 8.

OpenStudy (anonymous):

ur welcome , can i get a badge :)?

OpenStudy (anonymous):

thank you and how do i give badges?

OpenStudy (anonymous):

its next to the badge sign where it say "best Answer" press that